OpenStudy (derpisthewordorisit):

What is the product in lowest terms? 3/8*2/7 A.1/28 B.3/56 C.5/56 D.3/28

OpenStudy (derpisthewordorisit):

@CGGURUMANJUNATH

OpenStudy (jango_in_dtown):

3/8*2/7=(3*2)/(8*7)=3/28